OpenStudy (anonymous):

In a triangle one angle is 12 degrees more than another. The third angle is twice the smallest angle. Find the measure of the smallest angle

OpenStudy (anonymous):

let first angle be x, second be y and third be z y=x+12 z=2x x+y+z=180 x+x+12+2x=180 4x+12=180